OpenStudy (anonymous):

HELP I WILL GIVE MEDAL AND BECOME FAN For question 1-2, simplify the expression. 1. √3+4√3 A. 5√6 B. 5√3 C. 3√6 D. 3√3 2. 2√6+3√96 A. 14√6 B. 14√96 C. 5√96 D. 50√6

OpenStudy (anonymous):

1 - B 2 - A

OpenStudy (anonymous):

do u know how to solve them ??

OpenStudy (anonymous):

Can you please tell me how you got it? :) Thanks

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1- \[\sqrt{3} (1+4) = 5\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{6} (2 + 3 (\sqrt{16}) = \sqrt{6} (2+3(4)) = \sqrt{6} (14)\]

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

what ??